Give the general solution using determination of

Integrating Factors.

1. (x^2 + y^2 + 1)dx + x(x − 2y)dy = 0

2. y(4x + y)dx − 2((x^2 − y)dy = 0

3. y(y + 2x − 2)dx − 2(x + y)dy = 0

4. y(8x − 9y)dx − 2x(x − 3y)dy = 0

5. y(2x^2 − xy)dx − (x − y)dy = 0

6. 2(2y^2 + 5xy − 2y + 4)dx + x(2x + 2y − 1)dy = 0

7. (2x^2 + 3xy − 2y + 6x)dx + x(x + 2y − 1)dy = 0


1
Expert's answer
2021-05-07T09:00:08-0400

1. "(x^2 + y^2 + 1)dx + x(x \u2212 2y)dy = 0"


Integrating factor-

"I.F.=\dfrac{1}{Mdx+Ndy}=\dfrac{1}{\frac{x^3}{3}+xy^2+x}"

Given equation is of the form Mdx+Ndy, Integrate M w.r.t. x and integrate only those term which are

free from x-

The solution of the above equation is-


"=\dfrac{x^3}{3}+xy^2+x"

2. "y(4x + y)dx \u2212 2(x^2 \u2212 y)dy = 0"


Integrating factor-

"I.F.=\dfrac{1}{Mdx+Ndy}=\dfrac{1}{2yx^2+xy^2+y^2}"

Given equation is of the form Mdx+Ndy, Integrate M w.r.t. x and integrate only those term which are

free from x-

The solution of the above equation is-


"=2yx^2+xy+y^2"


3. "y(y + 2x \u2212 2)dx \u2212 2(x + y)dy = 0"



Integrating factor-

"I.F.=\dfrac{1}{Mdx+Ndy}=\dfrac{1}{x^2+xy^2-2xy-y^2}"

Given equation is of the form Mdx+Ndy, Integrate M w.r.t. x and integrate only those term which are

free from x-

The solution of the above equation is-


"=x^2+xy^2-2xy-y^2"



4. "y(8x \u2212 9y)dx \u2212 2x(x \u2212 3y)dy = 0" 4


Integrating factor-

"I.F.=\dfrac{1}{Mdx+Ndy}=\dfrac{1}{4x^2y-9y^2x+3xy^2}"

Given equation is of the form Mdx+Ndy, Integrate M w.r.t. x and integrate only those term which are

free from x-

The solution of the above equation is-


"=4x^2y-9xy^2+3xy^2"




5. "y(2x^2 \u2212 xy)dx \u2212 (x \u2212 y)dy = 0"


Integrating factor-

"I.F.=\dfrac{1}{Mdx+Ndy}=\dfrac{1}{\frac{2x^3y}{3}-\frac{y^2x^2}{2}+\frac{y^2}{2}}"

Given equation is of the form Mdx+Ndy, Integrate M w.r.t. x and integrate only those term which are

free from x-

The solution of the above equation is-


"=\dfrac{2x^3y}{3}-\dfrac{y^2x^2}{2}+\dfrac{y^2}{2}"


6. "2(2y^2 + 5xy \u2212 2y + 4)dx + x(2x + 2y \u2212 1)dy = 0"


Integrating factor-

"I.F.=\dfrac{1}{Mdx+Ndy}=\dfrac{1}{4y^2x+5x^2y-4yx+8x}"

Given equation is of the form Mdx+Ndy, Integrate M w.r.t. x and integrate only those term which are

free from x-

The solution of the above equation is-


"=4y^2x+5x^2y-4yx+8x"


7. "(2x^2 + 3xy \u2212 2y + 6x)dx + x(x + 2y \u2212 1)dy = 0"


Integrating factor-

"I.F.=\dfrac{1}{Mdx+Ndy}=\dfrac{1}{\frac{2x^3}{3}+\frac{3x^2y}{2}-2yx}"

Given equation is of the form Mdx+Ndy, Integrate M w.r.t. x and integrate only those term which are

free from x-

The solution of the above equation is-


"=\dfrac{2x^3}{3}+\dfrac{3x^2y}{2}-2y"